OpenStudy (anonymous):

How many grams of calcium phosphate can be produced when 89.3 grams of calcium chloride reacts with excess sodium phosphate? Unbalanced equation: CaCl2 + Na3PO4 → NaCl + Ca3(PO4)2 Show, or explain, all of your work along with the final answer.

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@nacholover

OpenStudy (anonymous):

Balanced equation: 3CaCl2+2Na3PO4→6NaCl+Ca3(PO4)2 Ca = 40.08g Cl = 35.45g Na = 11.99g P = 30.97g O = 16.00g It takes 3 moles of Calcium Chloride to produce 1 mole of Calcium Phosphate. 1 mole CaCl2(40.08g Ca+(2×35.45g Cl2))×89.3g CaCl21=0.805 moles 3 moles CaCl21 mole Ca3(PO4)2=0.805 moles CaCl2x moles Ca3(PO4)2 3.00x=0.805×1→x=0.8053.00→x=0.268 moles Ca3(PO4)2 89.3g of CaCl2 will produce 72.45g of Ca3(PO4)2

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

I have a few more but I will tag you when I post them!!

OpenStudy (anonymous):

just do it tommorow cus im about to go home from school

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

ok do you still need help/